# Physics-Aware Sparse Signal Recovery Through PDE-Governed Measurement Systems

### v1

@@ Abstract
This paper introduces a novel framework for physics-aware sparse signal recovery in measurement systems governed by partial differential equations (PDEs). Unlike conventional compressed sensing approaches that treat measurement systems as simple linear systems, our method explicitly incorporates the underlying physics through numerical PDE solvers and automatic differentiation (AD). We present physics-aware iterative shrinkage-thresholding algorithm (PA-ISTA), which combines the computational efficiency of ISTA with accurate physical modeling to achieve improved signal reconstruction. Using optical fiber channels as a concrete example, we demonstrate how the nonlinear Schrödinger equation (NLSE) can be integrated into the recovery process. Our approach leverages deep unfolding techniques for parameter optimization. Numerical experiments show that PA-ISTA significantly outperforms conventional recovery methods. While demonstrated on optical fiber systems, our framework provides a general methodology for physics-aware signal recovery that can be adapted to various PDE-governed measurement systems.
@@ I Introduction
Recent advances in computational techniques, particularly automatic differentiation (AD) [10], have opened new possibilities for incorporating physical models into signal processing frameworks. In this paper, we propose a general framework for physics-aware sparse signal recovery that incorporates PDEs describing the underlying physical phenomena. While our methodology is applicable to a wide range of physical systems governed by PDEs, we demonstrate its effectiveness using optical fiber channels [9] as a concrete and challenging example. The key innovation lies in our use of AD mechanism to compute gradients through a numerical PDE solver, enabling efficient optimization despite the complexity of the physical system. The proposed framework is general enough to be adapted to other physical systems where the underlying physics can be described by PDEs, such as heat conduction processes, wave propagation in elastic media, or fluid dynamics.
@@ II Preliminaries
II-A Sparse signal recovery problem
@@
where $\bm{y}\in\mathbb{R}^{m}$ $(m<n)$ represents the measurement vector, $\bm{A}\in\mathbb{R}^{m\times n}$ is the measurement matrix, and $\bm{w}\in\mathbb{R}^{m}$ denotes the measurement noise. In the sparse signal recovery problem, the objective is to recover the original signal $\bm{s}$ from the given measurement vector $\bm{y}$ with maximum possible accuracy. The observer has access only to $\bm{y}$ and $\bm{A}$. It is important to note that since $m<n$, without considering the sparsity of the original signal $\bm{s}$, this becomes an underdetermined problem with no unique solution.
@@
II-B ISTA for sparse signal recovery
@@
II-C Nonlinear Schrödinger Equation (NLSE)
@@
II-D State Split Fourier Method (SSFM)
The nonlinear Schrödinger equation involves both linear dispersion terms and nonlinear effects, making it challenging to solve directly. The split-step Fourier method (SSFM) [9] provides an efficient numerical approach by leveraging the distinct characteristics of these terms. Consider the nonlinear Schrödinger equation: ${\partial A}/{\partial z}=(\hat{D}+\hat{N})A,$ where $A$ is the complex envelope of the optical field, $\hat{D}$ represents the linear dispersion operator in the frequency domain, and $\hat{N}$ represents the nonlinear operator in the time domain.
@@ II-D State Split Fourier Method (SSFM)
This approach is particularly efficient because the dispersion operator $\hat{D}$ takes a simple multiplicative form in the frequency domain, while the nonlinear operator $\hat{N}$ is easily computed in the time domain. The FFT provides an efficient means to switch between these domains, making the overall method computationally practical.
III Sparse Signal Recovery Problem
@@ III-A Overview
Although our physics-aware sparse signal recovery framework is applicable to a wide range of partial differential equations, we demonstrate its effectiveness using the NLSE (7) as a concrete example. In our problem formulation, we consider a system where a sensing device at the fiber input ($z=0$) injects a waveform $U(t,0)$ consisting of several sparse pulses, which can be considered as the target sparse signals. As this signal propagates through the fiber, it is observed by a detector placed at the fiber output ($z=L$). Our goal is to reconstruct the target sparse signal from these output measurements, effectively recovering the input signal through the nonlinear fiber channel. As the optical signal propagates through the fiber, the waveform evolves according to the NLSE (7) with increasing $z$. The signal recovery problem can be stated as follows: given a noisy observation of the waveform $U(t,L)$ at position $z=L$, estimate the original sparse signal $U(t,0)$. In other words, we aim to detect the set of sparse pulses at the fiber input ($z=0$) using measurements obtained at the fiber output ($z=L$).
@@ III-B Details
$\phi(x)\equiv\exp\left(-{x^{2}}/{(2T_{0}^{2})}\right),$ (8)
@@
where each noise term $n_{i}$ follows a complex Gaussian distribution, i.e., $n_{i}\sim{\cal CN}(\bm{0},\sigma^{2})$. Figure 1 visualizes behavior of signal propagation through an optical fiber.
@@ IV-A Lasso-like formulation
where $f(\bm{s})\equiv(U(q_{1},L;\bm{s}),\ldots,U(q_{m},L;\bm{s}))$ represents the noiseless output of the nonlinear fiber channel at the sampling points. The first term measures the fidelity between the observed samples and the predicted output, while the second term promotes sparsity in the solution. By minimizing this objective function, we aim to obtain a sparse solution that is consistent with both the observed data and the underlying physics described by the NLSE.
@@ IV-B Physics-aware ISTA
where the objective function becomes non-convex when the fiber channel exhibits significant nonlinearity. Despite this non-convexity, we propose to adapt the ISTA framework to minimize $F(\bm{s})$. This leads to a complex-valued variant of the ISTA algorithm:
@@
We initialize the state vector using the zero-forcing solution obtained through the Backpropagation method (BP)  [^1] Note that this backpropgation method is not for computing the gradients. [13], a widely established technique for nonlinearity compensation in optical fiber communications. BP reconstructs the input signal by numerically solving the NLSE in reverse, with the signs of both dispersion and nonlinearity coefficients inverted. This approach effectively ”undoing” the channel effects by propagating the received signal backwards through a virtual fiber with inverse channel parameters.
In the ideal case of noiseless transmission, BP can achieve perfect signal recovery. However, in practical scenarios where noise is present, BP’s performance degrades significantly due to its inherent noise amplification characteristics.
The proposed algorithm exhibits a double-loop structure: an outer loop implementing the proximal gradient iterations, and an inner loop executing the SSFM. The computational complexity of the SSFM solver scales with the number of grid points, requiring $O(N_{t}N_{z})$ operations, where $N_{t}$ and $N_{z}$ represent the number of grids in $t$ and $z$ direction, respectively. When $\max\{n,m\}<\min\{N_{z},N_{t}\}$, the overall computational complexity of the algorithm becomes $O(UN_{t}N_{z})$. A practical trade-off exists between computational efficiency and solution accuracy. While using a coarser grid (smaller $N_{t}$ and $N_{z}$) reduces computational time, it may compromise the quality of the recovered signal. Therefore, careful selection of grid parameters is crucial for balancing computational complexity against recovery performance.
@@
1:  Set the initial state $\bm{x}^{(0)}:=BP(\bm{y})$.
2:  for $k:=0$ to $U-1$ do
3:     Let $b(t):=\sum_{i=1}^{n}x_{i}^{(k)}\phi(t-p_{i}).$
4:     Solve the NLSE (7) with the boundary condition $\hat{U}(t,0;\bm{x}^{(k)})=b(t)$ by using the SSFM solver.
5:     Generate reconstructed samples:
$r_{i}:=\hat{U}(q_{i},L;\bm{x}^{(k)}),\ i\in[m].$
6:     Let $r(\bm{x}^{(k)})\equiv(r_{1},r_{2},\ldots,r_{m})$.
7:     Compute the Wirtinger gradient by AD:
$\bm{g}:=\nabla_{\bm{x}^{(k)}}\|\bm{y}-\bm{r}(\bm{x}^{(k)})\|^{2}_{2}.$
8:     Execute the GD process: $\bm{z}^{(k)}:=\bm{x}^{(k)}-\eta^{(k)}\bm{g}.$
9:     Execute the shrinkage process: $\bm{x}^{(k+1)}:=T_{\theta^{(k)}}(\bm{z}^{(k)}).$
@@ IV-C Deep Unfolding (DU) for Parameter Tuning
To overcome this implementation challenge, we propose the store-and-replay method, which effectively decouples the two levels of AD by executing PA-ISTA iterations twice in a specific manner. The key idea is to separate the physics-based Wirtinger gradient computation from the DU-parameter optimization process. Algorithm 2 presents a pseudo code of the store-and-replay method. In the first pass (Store phase), we execute PA-ISTA iterations with AD enabled for the Wirtinger derivative. The second pass (Replay phase) then replicates the same PA-ISTA iterations but with a crucial difference; the gradients are retrieved from the storage. This two-pass approach successfully avoids nested AD while maintaining the mathematical equivalence to the original algorithm. The store-and-replay method enables us to apply DU for parameter tuning by effectively separating the physics-based gradient computation from the DU parameter optimization process.
@@
IV-D Estimation of NLSE Parameters
Our discussion of PA-ISTA has thus far assumed perfect knowledge of the NLSE parameters, specifically the dispersion coefficient $\beta_{2}$ and the nonlinearity coefficient $\gamma$. In practical scenarios, however, these parameters are often unknown and must be estimated from measurements. This parameter uncertainty presents an additional challenge for physics-aware signal recovery, as incorrect parameter values can significantly degrade the recovery performance of PA-ISTA.
When the assumed parameters in the SSFM solver deviate from their true values, the numerical solution of the NLSE no longer accurately represents the actual signal propagation, leading to incorrect gradient computations and subsequently poor recovery performance. This sensitivity highlights the importance of accurate parameter estimation for successful implementation of physics-aware recovery methods.
In oder to achieve accurate parameter estimation, we propose a parameter estimation method based on solving the following least square problem:
$\mbox{minimize }_{\beta_{2}\in\mathbb{R},\gamma\in\mathbb{R}}\|\bm{y}-\bm{r}(\bm{s})\|^{2}_{2},$ (20)
where $\bm{s}$ represents a known sparse pilot signal, and $\bm{y}$ is the corresponding noisy observation vector obtained at the detector. This formulation seeks to find the NLSE parameters that best explain the observed signal propagation for the known pilot signal. The optimization problem can be solved using gradient descent, where the gradients with respect to $\beta_{2}$ and $\gamma$ are efficiently computed through AD. This approach leverages the same computational framework used in PA-ISTA, making it naturally compatible with our physics-aware recovery scheme.
However, it is important to note that this minimization problem is inherently non-convex, introducing the possibility of the gradient descent process converging to undesirable stationary points. Consequently, the quality of the parameter estimates can be highly dependent on the initial values chosen for the optimization process. This sensitivity to initialization suggests the need for careful selection of starting points, possibly incorporating prior knowledge about typical parameter ranges in optical fiber systems.

### v2

@@ Abstract
This paper introduces a novel framework for physics-aware sparse signal recovery in measurement systems governed by partial differential equations (PDEs). Unlike conventional compressed sensing approaches that treat measurement systems as simple linear systems, our method explicitly incorporates the underlying physics through numerical PDE solvers and automatic differentiation (AD). We present physics-aware iterative shrinkage-thresholding algorithm (PA-ISTA), which combines the computational efficiency of ISTA with accurate physical modeling to achieve improved signal reconstruction. Using optical fiber channels as a concrete example, we demonstrate how the nonlinear Schrödinger equation (NLSE) can be integrated into the recovery process. Our approach leverages deep unfolding techniques for parameter optimization. Numerical experiments show that PA-ISTA significantly outperforms conventional recovery methods. While demonstrated on optical fiber systems, the proposed framework provides a general methodology for physics-aware signal recovery applicable to a wide range of various PDE-governed measurement systems.
@@ I Introduction
Recent advances in computational techniques, particularly automatic differentiation (AD) [15], have opened new possibilities for incorporating physical models into signal processing frameworks. In this paper, we propose a general framework for physics-aware sparse signal recovery that incorporates PDEs describing the underlying physical phenomena. While our methodology is applicable to a wide range of physical observation systems governed by PDEs, we demonstrate its effectiveness using optical fiber channels [14] as a concrete and challenging example.
In optical fiber systems, we integrate the computational efficiency of ISTA with numerical PDE solvers to achieve accurate signal reconstruction while faithfully capturing the underlying physics. The key innovation lies in our use of AD mechanism to compute gradients through a numerical PDE solver, thereby facilitating efficient optimization despite the complexity of the physical system. The proposed framework is general enough to be adapted to other physical systems where the underlying physics can be described by PDEs, such as heat conduction processes, wave propagation in elastic media, Maxwell’s equations for electromagnetic waves, or fluid dynamics. Figure 1 shows the scope for physics-aware sparse signal recovery presented in this paper.
Figure 1: Scope for physics-aware sparse signal recovery.
As a related work, the concept of physics-aware decoding that successfully uses automatic differentiation in a decoding process for binary linear codes has been presented in [27]. However, the extension to sparse signal recovery problems has not yet been discussed so far. The main contributions of this paper are as follows:
  • We propose a general framework for physics-aware sparse signal recovery that incorporates PDEs describing the underlying physical phenomena.
  • We demonstrate the effectiveness of the proposed framework using optical fiber channels as a concrete example.
  • Numerical experiments show that the proposed framework significantly outperforms conventional naive recovery methods.
The rest of this paper is organized as follows. In Section II, we review the fundamentals of compressed sensing and the ISTA algorithm. In Section III, we formulate the sparse signal recovery problem and introduce the proposed framework. In Section V, we present numerical experiments to demonstrate the effectiveness of the proposed framework. In Section VI, we draw conclusions and discuss future work.
We have released an implementation, including scripts to reproduce numerical results for sparse signal recovery, at https://github.com/wadayama/PA-ISTA.
@@ II Preliminaries
II-A Sparse signal recovery problem with linear measurement
@@
where $\bm{y}\in\mathbb{R}^{m}$ $(m<n)$ represents the measurement vector, $\bm{A}\in\mathbb{R}^{m\times n}$ is the measurement matrix, and $\bm{w}\in\mathbb{R}^{m}$ denotes the measurement noise. In the sparse signal recovery problem, the objective is to recover the original signal $\bm{s}$ from the given measurement vector $\bm{y}$ with maximum possible accuracy. The observer has access only to $\bm{y}$ and $\bm{A}$. It is important to note that since $m<n$, without considering the sparsity of the original signal $\bm{s}$, this becomes an underdetermined problem with no unique solution. This framework is widely applicable such as MRI for medical imaging, and grant-free wireless communication systems.
@@
II-B Related works
Various discrete-time algorithms for sparse signal recovery [3] have been proposed in the literature based on the Lasso framework [5, 4, 6]. Among them, the iterative shrinkage thresholding algorithm (ISTA) [7, 8] has emerged as one of the most widely used approaches for solving the Lasso problem. ISTA operates by iteratively applying two key processes: a linear estimation step followed by a shrinkage operation using a soft thresholding function. From an optimization perspective, ISTA can be interpreted as a proximal gradient descent method [17] that naturally arises from the Lasso formulation. The approximate message passing (AMP) algorithm [9, 10] represents another significant advance, offering substantially faster convergence compared to ISTA. A notable extension is the orthogonal AMP (OAMP) [11] developed by Ma and Ping, which broadens the applicability to diverse sensing matrix classes, particularly those exhibiting unitary invariance. Further theoretical progress within the Bayesian regime was achieved by Rangan et al. with vector AMP (VAMP) [12], provides rigorous theoretical guarantees through state evolution analysis.
Signal detection algorithms play a crucial role in receiver design for optical fiber communication systems. Digital back propagation (DBP) [18] is a well-established signal detection technique that leverages the NLSE solver in reverse way. The learnable variants of DBP introduced in [19, 20] represent pioneering contributions demonstrating the applicability of deep learning to signal processing in optical fiber communications.
II-C ISTA for sparse signal recovery
@@
II-D Nonlinear Schrödinger Equation (NLSE)
@@
II-E Split-State Fourier Method (SSFM)
The NLSE involves both linear dispersion terms and nonlinear effects, making it challenging to solve directly. The split-step Fourier method (SSFM) [14] provides an efficient numerical approach by leveraging the distinct characteristics of these terms. Consider the NLSE:
${\partial A}/{\partial z}=(\hat{D}+\hat{N})A,$ (8)
where $A$ is the complex envelope of the optical field, $\hat{D}$ represents the linear dispersion operator in the frequency domain, and $\hat{N}$ represents the nonlinear operator in the time domain.
@@ II-E Split-State Fourier Method (SSFM)
  5. 5.
    This process repeats for the next propagation step.
This approach is particularly efficient because the dispersion operator $\hat{D}$ takes a simple multiplicative form in the frequency domain, while the nonlinear operator $\hat{N}$ is easily computed in the time domain. The FFT provides an efficient means to switch between these domains, making the overall method computationally practical. The symmetrized version of SSFM, which applies half of the linear operation before and after the nonlinear step, achieves second-order accuracy with respect to the step size $\Delta z$. Specifically, it approximates the solution over a step $\Delta z$ as
$A(z+\Delta z)\approx\exp\left(\frac{\Delta z}{2}\hat{D}\right)\exp\left(\Delta z\hat{N}\right)\exp\left(\frac{\Delta z}{2}\hat{D}\right)A(z).$ (9)
This improved accuracy, combined with its computational efficiency, makes the symmetrized SSFM particularly suitable for our physics-aware recovery framework where both precision and computational tractability are essential. A pseudo-code of the SSFM is given in Algorithm 1 according to [14].
Algorithm 1 Split-Step Fourier Method (SSFM) for NLSE
0: Input waveform $U(t,0)$, fiber length $L$, spatial step size $\Delta z$, temporal grid $\{t_{1},t_{2},\ldots,t_{N_{t}}\}$, parameters $\beta_{2}$, $\gamma$
0: Output waveform $U(t,L)$
[…]
3:for $k=1$ to $N_{z}$ do
4:  $z\leftarrow k\cdot\Delta z$
5:  // Linear step (frequency domain) - first half
6:  $\tilde{U}\leftarrow\text{FFT}(U_{\text{current}})$
7:  for $i=1$ to $N_{t}$ do
8:   $\omega_{i}\leftarrow 2\pi(i-N_{t}/2)/(N_{t}\cdot\Delta t)$
9:   $\tilde{U}_{i}\leftarrow\tilde{U}_{i}\cdot\exp\left(\frac{i\beta_{2}\omega_{i}^{2}\Delta z}{4}\right)$
10:  end for
[…]
12:  // Nonlinear step (time domain)
13:  for $i=1$ to $N_{t}$ do
14:   $U_{\text{mid},i}\leftarrow U_{\text{mid},i}\cdot\exp(i\gamma|U_{\text{mid},i}|^{2}\Delta z)$
15:  end for
16:  // Linear step (frequency domain) - second half
17:  $\tilde{U}\leftarrow\text{FFT}(U_{\text{mid}})$
18:  for $i=1$ to $N_{t}$ do
19:   $\omega_{i}\leftarrow 2\pi(i-N_{t}/2)/(N_{t}\cdot\Delta t)$
20:   $\tilde{U}_{i}\leftarrow\tilde{U}_{i}\cdot\exp\left(\frac{i\beta_{2}\omega_{i}^{2}\Delta z}{4}\right)$
21:  end for
[…]
23:end for
24:
25:return $U_{\text{current}}$
III Sparse Signal Recovery Problem for PDE-governed Measurement Systems
@@ III-A Overview
Although our physics-aware sparse signal recovery framework is applicable to a wide range of partial differential equations, we demonstrate its effectiveness using the NLSE (7) as a concrete example. In our problem formulation, we consider a system where a sensing device at the fiber input ($z=0$) injects a waveform $U(t,0)$ consisting of several sparse pulses, which can be considered as the target sparse signals. As this signal propagates through the fiber, it is observed by a detector placed at the fiber output ($z=L$). Our goal is to reconstruct the target sparse signal from these output measurements, effectively recovering the input signal through the nonlinear fiber channel. Specifically, as the optical signal propagates through the fiber, the waveform evolves according to the NLSE (7) with increasing $z$. The signal recovery problem can be stated as follows: given a noisy observation of the waveform $U(t,L)$ at position $z=L$, estimate the original sparse signal $U(t,0)$. In other words, we aim to detect the set of sparse pulses at the fiber input ($z=0$) using measurements obtained at the fiber output ($z=L$).
@@ III-B Details
$\phi(x)\equiv\exp\left(-\frac{x^{2}}{2T_{0}^{2}}\right),$ (10)
@@
where each noise term $n_{i}$ follows a complex Gaussian distribution, i.e., $n_{i}\sim{\cal CN}(\bm{0},\sigma^{2})$ where $\sigma^{2}$ denotes noise variance. Figure 2 visualizes behavior of signal propagation through an optical fiber.
@@ IV-A Lasso-like formulation
where
$f(\bm{s})\equiv(U(q_{1},L;\bm{s}),\ldots,U(q_{m},L;\bm{s}))$ (14)
represents the noiseless output of the nonlinear fiber channel at the sampling points. The first term measures the fidelity between the observed samples and the predicted output, while the second term promotes sparsity in the solution. By minimizing this objective function, we aim to obtain a sparse solution that is consistent with both the observed data and the underlying physics described by the NLSE.
@@ IV-B Physics-aware ISTA
where the objective function becomes non-convex when the optical fiber channel exhibits significant nonlinearity. Despite this non-convexity, we propose to adapt the ISTA framework to minimize $F(\bm{s})$. This leads to a complex-valued variant of the ISTA algorithm:
@@
The signal recovery process of PA-ISTA is also depicted in Figure 3. From the estimated signal $\bm{x}^{(k)}$, the initial waveform is generated as a boundary condition of the NLSE. The SSFM solver then propagates the waveform through the fiber, yielding the predicted output $\hat{f}(\bm{x}^{(k)})$. The difference between the predicted output and the observed samples is then used to update the state vector $\bm{x}^{(k+1)}$. This process is repeated until the state vector converges to a sparse solution or the number of iterations reaches the predefined maximum value.
Figure 3: Signal recovery process of PA-ISTA.
We initialize the state vector using the zero-forcing solution obtained through the digital back propagation(DBP) [18], a widely established technique for nonlinearity compensation in optical fiber communications. DBP reconstructs the input signal by numerically solving the NLSE in reverse, with the signs of both dispersion and nonlinearity coefficients inverted. This approach effectively ”undoing” the channel effects by propagating the received signal backwards through a virtual fiber with inverse channel parameters.
In the ideal case of noiseless transmission, DBP can achieve almost perfect signal recovery. However, in practical scenarios where noise is present, DBP’s performance degrades significantly due to its inherent noise amplification characteristics.
@@
0: Observations $\mathbf{y}$, observation position $L$, NLSE parameters $\beta_{2}$, $\gamma$, step size parameters $\bm{\eta}$, shrinkage parameters $\bm{\theta}$, iterations $U$
0: Estimated signal $\hat{\mathbf{s}}$
[…]
2:for $k=0$ to U1U-1 do
3:  $b(t)\leftarrow\sum_{i=1}^{n}x_{i}^{(k)}\phi(t-p_{i})$ {Construct input waveform}
4:  Solve the NLSE with boundary condition
$\hat{U}(t,0;\mathbf{x}^{(k)})=b(t)$
using SSFM solver
5:  $r_{i}\leftarrow\hat{U}(q_{i},L;\mathbf{x}^{(k)}),\,i\in[m]$ {Generate reconstructed samples}
6:  $\mathbf{r}(\mathbf{x}^{(k)})\leftarrow(r_{1},r_{2},\ldots,r_{m})$ {Construct measurement vector}
7:  $\mathbf{g}\leftarrow\nabla_{\mathbf{x}^{(k)}}\|\mathbf{y}-\mathbf{r}(\mathbf{x}^{(k)})\|^{2}_{2}$ {Compute Wirtinger gradient by AD}
8:  $\mathbf{z}^{(k)}\leftarrow\mathbf{x}^{(k)}-\eta^{(k)}\mathbf{g}$ {Gradient descent step}
9:  $\mathbf{x}^{(k+1)}\leftarrow\mathcal{T}_{\theta^{(k)}}(\mathbf{z}^{(k)})$ {Shrinkage step}
@@ IV-C1 Nested structure of gradient computation
Figure 4: Nested structure of gradient computation (A) and store-and-replay method (B).
IV-C2 Store-and-Replay Method
To overcome this implementation challenge, we propose the store-and-replay method, which effectively decouples the two levels of AD by executing PA-ISTA iterations twice in a specific manner. The key idea is to separate the physics-based Wirtinger gradient computation from the DU-parameter optimization process. Algorithm 3 presents a pseudo code of the store-and-replay method. In the first pass (Store phase), we execute PA-ISTA iterations with AD enabled for the Wirtinger derivative. The second pass (Replay phase) then replicates the same PA-ISTA iterations but with a crucial difference; the gradients are retrieved from the storage. This two-pass approach successfully avoids nested AD while maintaining the mathematical equivalence to the original algorithm. The store-and-replay method enables us to apply DU for parameter tuning by effectively separating the physics-based gradient computation from the DU parameter optimization process. The details of the DU-training process for PA-ISTA is given in Algorithm 4.
@@ IV-C2 Store-and-Replay Method
Algorithm 4 Training procedure for PA-ISTA with deep unfolding
0: Target signal $\mathbf{s}_{0}$, observations $\mathbf{y}$, observation position $z_{obs}$, learnable parameters $\bm{\eta}$, $\bm{\theta}$, optimizer optopt, iterations $U$, NLSE parameters $\beta_{2}$, $\gamma$
0: Updated parameters $\bm{\eta}$, $\bm{\theta}$
[…]
2: // Store Phase - Compute and store gradients
[…]
4:for $k=1$ to $U$ do
5:  $f(\mathbf{x})=\|\mathbf{y}-\text{SSFM}(\mathbf{x},z_{obs},\beta_{2},\gamma)\|^{2}_{2}$ {Define loss function}
6:  $\nabla\mathbf{s}\leftarrow\text{WirtingerGrad}(f,\mathbf{s})$ {Compute gradient via AD}
7:  $\nabla\text{History}[:,k]\leftarrow\nabla\mathbf{s}$ {Store gradient}
8:  $\mathbf{s}\leftarrow\mathbf{s}-|\eta_{k}|\cdot\nabla\mathbf{s}$ {Gradient step}
9:  $\mathbf{s}\leftarrow\mathcal{T}_{\theta_{k}}(\mathbf{s})$ {Shrinkage step}
10:end for
11: // Replay Phase - Update parameters using stored gradients
[…]
14: Compute gradients $\nabla_{\bm{\eta}}\mathcal{L}$, $\nabla_{\bm{\theta}}\mathcal{L}$ via backpropagation through:
15:for $k=1$ to $U$ do
16:  $\mathbf{s}\leftarrow\mathbf{s}-|\eta_{k}|\cdot\nabla\text{History}[:,k]$ {Use stored gradients}
17:  $\mathbf{s}\leftarrow\mathcal{T}_{\theta_{k}}(\mathbf{s})$ {Apply shrinkage}
18:end for
19: Update $\bm{\eta}$, $\bm{\theta}$ using an optimizer (e.g., Adam) with gradients $\nabla_{\bm{\eta}}\mathcal{L}$, $\nabla_{\bm{\theta}}\mathcal{L}$
20:
21:return Updated parameters $\bm{\eta}$, $\bm{\theta}$
IV-D Convergence and Computational Complexity
IV-D1 Convergence
Consider the objective
$F(\bm{s})=\tfrac{1}{2}\|\bm{y}-\hat{f}(\bm{s})\|_{2}^{2}+\lambda\|\bm{s}\|_{1},$
where $\hat{f}$ is the differentiable forward map obtained by the SSFM-based solver and AD (Wirtinger calculus). Although $F$ can be nonconvex due to the nonlinear forward operator, the data-fidelity gradient $\nabla d(\bm{s})\equiv\nabla\tfrac{1}{2}\|\bm{y}-\hat{f}(\bm{s})\|_{2}^{2}$ is locally Lipschitz under a mild regularity assumption on $\hat{f}$ (bounded Jacobian on the level set of interest). Let $\{\bm{s}^{k}\}$ be generated by the PA-ISTA iteration
$\bm{s}^{k+1}=\mathrm{prox}_{\eta_{k}\lambda\|\cdot\|_{1}}\!\big(\bm{s}^{k}-\eta_{k}\nabla d(\bm{s}^{k})\big)$ (23)
with either (i) backtracking to ensure a sufficient decrease condition, or (ii) fixed steps $\eta_{k}\in(0,\bar{\eta}]$ with $\bar{\eta}<1/L$ where $L$ is a local Lipschitz constant of $\nabla d$. For backtracking, we can shrink $\eta_{k}\leftarrow\tau\,\eta_{k}$ with $\tau\in(0,1)$ until $F(\bm{s}^{k+1})\leq F(\bm{s}^{k})-\frac{\sigma}{2\eta_{k}}\,\|\bm{s}^{k+1}-\bm{s}^{k}\|_{2}^{2}$ holds for some fixed $\sigma\in(0,1)$. Then every limit point of $\{\bm{s}^{k}\}$ is a first-order critical point of $F$ (proximal-stationary point), and $F(\bm{s}^{k})$ is nonincreasing. This is the standard proximal-gradient convergence behavior for smooth + nonsmooth composite objectives and remains valid in the present setting because the nonsmooth part is convex ($\ell_{1}$) and the gradient of the smooth part is (locally) Lipschitz. Deep unfolding keeps the iteration form but removes a priori guarantees unless the learned stepsizes obey a safeguard (e.g., $\eta_{k}\leq(1-\epsilon)/L$). With either such a safeguard or backtracking ensuring sufficient decrease, $F(\bm{s}^{k})$ is nonincreasing and every limit point is proximal-stationary.
IV-D2 Computational complexity
The proposed algorithm exhibits a double-loop structure: an outer loop implementing the proximal gradient iterations, and an inner loop executing the SSFM. The computational complexity of the SSFM solver scales with the number of grid points, requiring $O(N_{t}N_{z})$ operations, where $N_{t}$ and $N_{z}$ represent the number of grids in $t$ and $z$ direction, respectively. The overall computational complexity of the algorithm becomes $O(UN_{t}N_{z}\log N_{t})$ where $N_{t}\log N_{t}$ comes from the complexity of FFT. A practical trade-off exists between computational efficiency and solution accuracy. While using a coarser grid (smaller $N_{t}$ and $N_{z}$) reduces computational time, it may compromise the quality of the recovered signal. Therefore, careful selection of grid parameters is crucial for balancing computational complexity against recovery performance.
The memory use is dominated by the SSFM state and AD tapes. The store-and-replay strategy decouples nested AD, so peak memory is $\mathcal{O}(N_{t})$ per SSFM slice plus $\mathcal{O}(nU)$ for the stored Wirtinger gradients, while runtime remains linear in $U,N_{t},N_{z}$. Therefore, PA-ISTA exhibits near-linear scaling in grid resolution and iteration count, and its wall-clock can be traded off against accuracy through the choice of $(N_{t},N_{z})$ and $U$, as documented in the method description.
IV-E QPSK detection with PA-ISTA
The shrinkage function is a crucial component that incorporates prior information about the source signal, such as sparsity, into the recovery process. Although soft thresholding is commonly used for sparse signal recovery, one of PA-ISTA’s key strengths is its ability to handle diverse source signals through appropriate modifications to the shrinkage function. By simply adjusting this function within the PA-ISTA framework, we can effectively leverage different types of prior information about the characteristics of the source signal.
We here consider an application of PA-ISTA for quadrature phase-shift keying (QPSK) detection. In this scenario, the input signal is assumed to be a QPSK signal, and the observation is the received signal at the detector. The goal is to estimate the transmitted signal from the observation. This is a simpliefied scenario of the QPSK detection problem, and the PA-ISTA can be applied to this problem. The prior knowledge of the QPSK signal can be reflected by the shrinkage function. We can expect that the shrinkage function is matched to the QPSK signal provides better detection performance compared with the case of a naive DBP-based recovery.
Assume that the QPSK signal constellation is given by
$\mathcal{Q}\equiv\left\{1+i,-1+i,-1-i,1-i\right\},$ (24)
i.e., a transmitted signal is chosen from the set $\mathcal{Q}$ with equal probability. The initial waveform is generated from the QPSK signal with equal probability:
$U(\tau,0)=\sum_{i=1}^{n}s_{i}\phi(\tau-p_{i}),$ (25)
where $s_{i}\in\mathcal{Q}$ and $p_{i}$ is the position of the $i$-th element. The rest of the observation process is the same as the case of the sparse signal recovery discussed in the previous section.
In this case, the process of the PA-ISTA is modified as
$\bm{x}^{(k+1)}$ $=S(\bm{z}^{(k)},\theta^{(k)}),\quad k=0,1,\dots,$ (27)
where $S:\mathbb{C}\times\mathbb{R}\rightarrow\mathbb{C}$ is the shrinkage function matched to the QPSK signal. The shrinkage function is given by
$S(y,\lambda)$ $\equiv\tanh(\lambda\Re(y))+i\tanh(\lambda\Im(y)),$ (28)
where $\Re(y)$ and $\Im(y)$ denote the real and imaginary parts of $y$, respectively. The parameter $\lambda$ is a scaling factor that controls the strength of the shrinkage. The final estimate is given by
$\hat{\bm{s}}=\operatorname{proj}_{\mathcal{Q}}(\bm{x}^{(U)}),$ (29)
where $\operatorname{proj}_{\mathcal{Q}}(a+ib)=\operatorname{sign}(a)+i\,\operatorname{sign}(b)$.
The above detection process can be regarded as a projected gradient descent method with the soft projection operator given by the set $\mathcal{Q}$. In the context of MIMO signal detection, such a detection process is discussed in [26]. In optical fiber communications, signal detection is a key process for the receiver design. An iterative algorithm like PA-ISTA can be used for the signal detection to improve the detection performance.